OpenStudy (anonymous):

using the law of exponents. 4 to the 6th power multiplied by 4 to the negative 4 power.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is it multiplication or subtraction? just trying to clarify.

OpenStudy (saifoo.khan):

\[\huge 4^6 x 4^{-4}\]

OpenStudy (anonymous):

your problem says multiplication

OpenStudy (anonymous):

its multiplication sorry

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

If you multiply powers of the same base you add their exponents: \[\large a^b \cdot a^c = a^{b+c}\]

OpenStudy (anonymous):

So: \[4^6\cdot 4^{-4} = 4^{6+(-4)}\]

OpenStudy (saifoo.khan):

\[\huge 4^{6+(-4)} \] 4^{6-4} 4^2 16